solve equation by factoring: (y-3)(y+2)= 0
Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

(y-3)(y+2)= 0 It is factored! So, use the zero product property.
y-3=0 or y=3
y+2=0 or y=-2

Answer by stanbon(75887)

solve equation by factoring: (y-3)(y+2)= 0
Since the product is zero one of these factors must be zero;
so, y-3=0 or y+2=0
and, y=3 or y=-2
Cheers,
Stan H.